# A Nuclear Interferometer for Ultra-Light Dark Matter Detection

### v2

@@ V Conclusions
With an extraordinarily broad landscape of viable candidates which give rise to a diverse range of phenomenological signatures, new ideas for the detection of DM are required in order to maximise experimental coverage of the vast available parameter space. In this work we propose two possible realisations of a single-photon interferometry experiment based on the nuclear clock transition in 229Th as a probe of ULDM with scalar couplings to the SM, making use of single-ions and clouds of neutral atoms respectively. Although working with either source type presents a unique set of experimental challenges, we find that the enhanced sensitivity of the nuclear clock transition to the variation of fundamental constants offers the means to access large regions of uncharted parameter space over a wide range of frequencies providing experimental challenges can be met in the future.
@@ Acknowledgements.
This work is supported by the CERN Quantum Technology Initiative. We are incredibly grateful to Leondardo Badurina and Thomas Hird for discussions regarding the challenges of using sources with short excited state lifetime in atom interferometry and to Sebastian Ellis for raising the issue of magnetic field noise after submission of V1, reflected in the extended discussion in this V2. We also thank Dennis Schlippert, Marianna Safranova, Masha Baryakhtar, Gilad Perez and Ekkehard Peik for invaluable feedback on early versions of this manuscript, Michael Doser for discussions on the nuclear clock and John Carlton for insight into the AION experiment and gravitational gradient noise. We thank Jeremiah Mitchell for discussions regarding the treatment of noise sources and Nathaniel Sherrill and Abhishek Banerjee for details on the analyses of atomic clock comparison experiments. HB acknowledges partial support from the STFC HEP Theory Consolidated grants ST/T000694/1 and ST/X000664/1 and thanks other members of the Cambridge Pheno Working Group for useful discussions. EF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967, and via the SFB/CRC 1227 (DQ-mat) – Project-ID 274200144 – of the DFG.
@@ Appendix A Sensitivity to ULDM Couplings
Here we briefly outline how the experimental sensitivity given in Eq. 18 follows from Eq. 15, the signal amplitude for a single measurement.
The experimental sensitivity takes into account a total of $N$ phase measurements made continuously at a regular sampling rate $1/\Delta t$ over the experiment run time $T_{\textnormal{int}}$. We denote these measurements as $\{\Phi_{m}\}$, where $m$ labels the phase recorded from the interferometer sequence starting at time $m\Delta t$.
A discrete set of measurements made at different times can be conveniently characterised by its power spectral density (PSD), which contains information on both the amplitude and the frequency spread of the data. The PSD can be constructed via the Fourier transform of the measured signal
$\tilde{\Phi}_{k}=\sum_{m=0}^{N-1}\Phi_{m}\textnormal{exp}\left(-\frac{2\pi imk}{N}\right)~{},$ (36)
where $k$ takes integer values between 0 and $N-1$.
Using this definition, the PSD, $S_{k}$, can be expressed as
$S_{k}=\frac{(\Delta t)^{2}}{T_{\textnormal{int}}}\left|\tilde{\Phi}_{k}\right|^{2}~{}.$ (37)
Defining the angular frequency as $\omega=2\pi k/T_{\textnormal{int}}$, in the limit of shot-noise domination, the SNR at a frequency $\omega_{0}$ is then given by
@@
where $S_{\textnormal{s}}(\omega)$ denotes the PSD of the ULDM induced signal and $S_{\textnormal{n}}(\omega)$ denotes the shot-noise PSD as given by Eq. 19.
In the limit $T_{\textnormal{int}}<\tau_{c}$, the ULDM signal is dominated by a single frequency such that the signal PSD takes the form
$S_{s}(\omega)\sim\begin{cases}T_{\textnormal{int}}|\overline{\Phi}_{s}|^{2},&\omega=\omega_{0}\\ 0,&\omega\neq\omega_{0}\end{cases}~{}.$ (39)
The experimental sensitivity to these couplings in this regime is thus
@@
If instead $T_{\textnormal{int}}>\tau_{c}$, the signal PSD has a finite profile in frequency space determined by the DM velocity distribution. In this instance, the experimental sensitivity can be computed using Bartlett’s method [92, 159, 160], leading to
@@
Combining these two results leads directly to Eq. 18.

### v3

@@ V Conclusions
With an extraordinarily broad landscape of viable candidates which give rise to a diverse range of phenomenological signatures, new ideas for the detection of DM are required in order to maximise experimental coverage of the vast available parameter space. In this work we propose two possible realisations of a single-photon interferometry experiment based on the nuclear clock transition in 229Th as a probe of ULDM with scalar couplings to the SM, making use of single-ions and clouds of neutral atoms respectively. Although working with either source type presents a unique set of experimental challenges, we find that the enhanced sensitivity of the nuclear clock transition to the variation of fundamental constants offers the means to access large regions of uncharted parameter space over a wide range of frequencies, providing experimental challenges can be met in the future.
@@ Acknowledgements.
This work is supported by the CERN Quantum Technology Initiative. We are incredibly grateful to Leondardo Badurina and Thomas Hird for discussions regarding the challenges of using sources with short excited state lifetime in atom interferometry and to Sebastian Ellis for raising the issue of magnetic field noise after submission of V1, reflected in the extended discussion in V2. We also thank Dennis Schlippert, Marianna Safronova, Masha Baryakhtar, Gilad Perez and Ekkehard Peik for invaluable feedback on early versions of this manuscript, Michael Doser for discussions on the nuclear clock and John Carlton for insight into the AION experiment and gravitational gradient noise. We thank Jeremiah Mitchell for discussions regarding the treatment of noise sources and Nathaniel Sherrill and Abhishek Banerjee for details on the analyses of atomic clock comparison experiments. HB acknowledges partial support from the STFC HEP Theory Consolidated grants ST/T000694/1 and ST/X000664/1 and thanks other members of the Cambridge Pheno Working Group for useful discussions. EF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967, and via the SFB/CRC 1227 (DQ-mat) – Project-ID 274200144 – of the DFG. This work has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 491245950.
@@ Appendix A Sensitivity to ULDM Couplings
Here we briefly outline how the experimental sensitivity given in Eq. 18 follows from Eq. 15, the signal amplitude for a single measurement. The experimental sensitivity takes into account a total of $N$ phase measurements made continuously at a regular sampling rate $1/\Delta t$ over the total experiment run time $T_{\textnormal{int}}$. We denote this series of measurements as $\{\Phi_{m}\}$, where $m$ labels the phase recorded from the interferometer sequence starting at time $m\Delta t$.
Such a discrete set of sequential measurements can be conveniently characterised by its power spectral density (PSD), which contains information on both the amplitude and the frequency spread of the data. It is formally constructed from the discrete Fourier transform of the sequence of measurements which itself can be calculated according to
$\tilde{\Phi}^{k}=\sum_{m=0}^{N-1}\Phi_{m}\textnormal{exp}\left(-\frac{2\pi imk}{N}\right)~.$ (36)
Here $k$ takes integer values between 0 and N1N-1 and labels the discrete (angular) frequencies $\omega=2\pi k/T_{\textnormal{int}}$ with resolution $\Delta\omega=2\pi/T_{\rm int}$.
Using this definition, the signal PSD, $S_{s}^{k}$, is
$S^{k}_{s}=\frac{(\Delta t)^{2}}{T_{\textnormal{int}}}\left|\tilde{\Phi}^{k}\right|^{2}~.$ (37)
In the limit that $T_{\textnormal{int}}<\tau_{c}$, the ULDM signal is contained within a single frequency bin. In this instance the signal PSD takes the form
$S_{s}(\omega)\sim\begin{cases}T_{\textnormal{int}}|\overline{\Phi}_{s}|^{2},&\omega=\omega_{0}\\ 0,&\omega\neq\omega_{0}\end{cases}~,$ (38)
where for convenience we have rewritten the PSD as a (continuous) function of angular frequency, $\omega=2\pi k/T_{\textnormal{int}}$.
The SNR can then be constructed as the ratio of the signal and noise PSDs evaluated at $\omega_{0}$
@@
where, assuming the experiment is shot-noise dominated, the noise power spectrum $S_{\textnormal{n}}$ is flat, and given by Eq. 19 and Eq. 33 for the ionic and atomic realisations respectively.
The sensitivity to the couplings of interest in this regime is thus
@@
If instead $T_{\textnormal{int}}>\tau_{c}$, the signal PSD has a finite profile in frequency space deriving from the DM velocity distribution. In this instance the experimental sensitivity is most rigorously extracted via a likelihood analysis as in e.g. Refs. [166, 167]. Alternatively one may apply Bartlett’s method [96, 168, 169].
The resulting sensitivity in this case is
@@
Combining this result with Eq. 40 leads directly to Eq. 18.
@@ Appendix C AION Prospective Sensitivity
Appendix D Sensitivity to the (light) QCD axion
For the purpose of illustrating the concept of nuclear interferometry we have thus far focused on the sensitivity to ULDM with linear scalar couplings to the SM as defined by the Lagrangian density in Eq. 1. In this appendix, we demonstrate that the proposed experiment is not exclusive to this particular model and can be used to probe other new physics scenarios which induce the apparent oscillation of fundamental constants. One particularly straightforward example of such a model (which we will not consider further here) is that of spin-0 scalar ULDM with quadratic couplings to the SM see e.g. Refs. [86, 87].
A further instance is that of QCD axion DM, which at phenomenologically viable masses, also behaves as a persistent classic wave, albeit with pseudo-scalar couplings to the SM. It was recently pointed out that QCD axion DM generates periodic temporal variations of atomic transition energies at quadratic order in the axion field by inducing oscillations in hadronic masses and other nuclear quantities [^20] In this work we focus on the oscillations in the 229Th transition energy that result from the quadratic couplings of the axion to hadrons that emerge below the QCD scale from the defining axion-gluon coupling. Oscillations can also be induced by the quadratic coupling of the QCD axion to photons which arises at 1-loop level due to threshold corrections [163, 164]. These are loop-suppressed relative to the hadronic couplings and we thus do not consider these further here. [162].
The defining feature of QCD axion models is an anomalous coupling between the pseudo-scalar axion field, $a$, and the SM gluon fields
$\mathcal{L}\in\frac{g_{s}^{2}}{32\pi^{2}}\frac{a}{f_{a}}G^{\mu\nu}\widetilde{G}_{\mu\nu}~,$ (45)
where $f_{a}$ is the axion decay constant, $g_{s}$ the strong coupling constant and $G$, ($\widetilde{G}$) the (dual) gluon field strength tensor. Below the QCD confinement scale, this interaction induces a quadratic coupling of the axion to hadronic states. In particular, the pion mass obtains a dependence on the axion field according to
$m_{\pi}^{2}(\theta_{\textnormal{eff}})=\frac{\Lambda^{3}_{\textnormal{QCD}}}{f_{\pi}^{2}}\sqrt{m_{u}^{2}+m_{d}^{2}+2m_{u}m_{d}\cos(\theta_{\textnormal{eff}})}~,$ (46)
where, for a QCD DM axion with mass $m_{a}$,
$\theta_{\textnormal{eff}}=\frac{(a-\langle a\rangle)}{f_{a}}=\frac{\sqrt{2\rho_{DM}}}{m_{a}f_{a}}\cos(m_{a}t)~.$ (47)
Given that it behaves as a coherently oscillating classical wave, QCD axion DM thus induces an oscillatory contribution to the pion mass at quadratic order in the axion field:
$\frac{\Delta m_{\pi}^{2}}{m_{\pi}^{2}}\sim\frac{m_{u}m_{d}\theta_{\textnormal{eff}}^{2}(t)}{2(m_{u}+m_{d})^{2}}~.$ (48)
This in turn causes the 229Th nuclear transition energy to oscillate as [162]
$\frac{\Delta\omega_{N}}{\omega_{N}}\sim(2\times 10^{5})\frac{\Delta m_{\pi}^{2}}{m_{\pi}^{2}}~.$ (49)
Combining Eq. 47 and Eq. 48, one can read off the amplitude of $\Delta\omega_{N}$ (defined in analogy to Eq. 12) to be
$\overline{\Delta\omega_{N}}=\omega_{N}(2\times 10^{5})\left(\frac{m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\right)\left(\frac{\rho_{DM}}{m_{a}^{2}f_{a}^{2}}\right)~.$ (50)
When the transition energy depends quadratically on the oscillating field as we have here, the single-shot signal amplitude (c.f. Eq. 15) for a gradiometer set-up takes the modified form
$\begin{split}\overline{\Phi}_{s,a}=2\frac{\overline{\Delta\omega_{N}}}{m_{a}}\frac{\Delta z}{L}\sin\left[m_{a}(T-(n-1)L)\right]\\ \times\sin\left[m_{a}nL\right]\sin\left[m_{a}T\right]~.\end{split}$ (51)
Defining, in analogy to Eq. 16,
$\overline{\Phi}_{s,a}=\frac{\overline{\Phi}_{R,a}}{f_{a}^{2}}~,$ (52)
it follows that the minimum value of the inverse axion decay constant that could be detected at a given SNR following a measurement campaign of duration $T_{\textnormal{int}}$ is
$\frac{1}{f_{a}^{*}}\simeq\sqrt{\frac{\sqrt{\textnormal{SNR}}}{\overline{\Phi}_{R,a}}\times\sqrt{\frac{S_{\textnormal{n}}}{T_{\textnormal{int}}}\times\textnormal{max}\left(1,\sqrt{\frac{T_{\textnormal{int}}}{\tau_{c}}}\right)}}~.$ (53)
where, as before, $S_{n}$ denotes the flat shot noise PSD as given in Eq. 19.
Figure 8: Shot-noise limited projected sensitivity of the nuclear interferometer configurations deploying neutral atoms detailed in Tab. 2 to QCD axion DM via its quadratic couplings to hadrons. The curves show the minimum value of $1/f_{a}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$ s and assuming that photo-ionisation has been successfully suppressed. The dotted portion of the curves indicates where the sensitivity may be adversely impacted by GGN. Also shown are existing bounds from astrophysics and other experiments as detailed in the main text, in addition to the forecast for nuclear-atomic clock frequency comparisons. The red QCD axion band indicates the expected parameter space coverage of standard QCD axion models in which the axion mass is generated by the anomalous coupling to quarks.
Figure 9: Shot-noise limited projected sensitivity of the single-ion nuclear interferometer configurations detailed in Tab. 1 to QCD axion DM via its quadratic couplings to hadrons. The curves show the minimum value of $1/f_{a}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$s. The dotted portion of the curve indicates where the sensitivity may be adversely impacted by magnetic field noise. Also shown are existing bounds from astrophysics and other experiments as detailed in the main text, in addition to the forecast for nuclear-atomic clock frequency comparisons. The red QCD axion band indicates the expected parameter space coverage of standard QCD axion models in which the axion mass is generated by the anomalous coupling to quarks.
[…]
For comparison we show the limits from an oscillating neutron EDM  [174] and beam based EDM experiment [175], in addition to a Rb/Cs hyperfine frequency clock comparison [152] and a frequency comparison of a hydrogen maser and a silicon cavity (H/Si) [154], using the limits derived in [162]. The limits from a frequency comparison of the E3/E2 transitions in Yb+ [157] and MICROSCOPE [176] are also displayed. In addition to laboratory experiments, we also show constraints from Supernova 1987 [177], white dwarfs [178], the binary neutron star gravitational wave event GW170817 [179] and big bang nucleosynthesis (BBN) [180]. Also displayed are bounds from binary pulsars [181] and a neutrino line from the solar core [181], using the corrected limits presented in Ref. [182] in both instances. Superradiance constraints from both stellar mass black holes and supermassive black holes (SMBH) are also shown  [183, 184, 185, 186, 187]. Although there are a variety of other proposals forecast to place constraints on this parameter space e.g. Refs. [188, 189, 190, 191], other than the projection for the frequency comparison of a nuclear and atomic clock [162], we do not display these projections here to avoid overcrowding.
Parameter space outside the red ‘QCD axion’ band corresponds to models which solve the strong CP problem, but have a mass that differs (accounting for theoretical uncertainty) from that generated by the anomalous coupling to gluons (Eq. 45) which defines $f_{a}$. Populating these regions generally requires significant model building and often suffers from naturalness problems.
From these figures, it is evident that the proposed realisations of the nuclear interferometer presented here are only able to probe ‘light’ QCD axions, whose masses are substantially less than that predicted from their defining QCD coupling. One example of a ‘natural’ model in which this feat is achieved is presented in Ref. [192] and subsequently developed phenomenologically in Ref. [182].
Whilst both the atomic and ionic nuclear interferometer configurations are able to reach substantial regions of parameter space not probed by existing laboratory experiments (and are on a par with forecasts from nuclear-atomic clock comparisons), this space has already been ruled out by astrophysical considerations. A nuclear interferometer could nonetheless provide a useful independent verification of these limits in a controlled laboratory environment.